# Robin3D[Uncaptioned image]  : Improving 3D Large Language Model via Robust Instruction Tuning

### v1

@@ 4.2 Diverse Data Generation
We first collect large scale data from different tasks outside the benchmark dataset. Specifically, given a 3D scene, we collect question-answering pairs from the following tasks: 1) Category Question-Answering task from Huang et al. [2023a], where the model is asked to answer the category of a specified object. 2) Nr3D Captioning task from Huang et al. [2023a], where the model is asked to caption the spatial relationship of a specified object to its neighboor. The ground truth is constructed from Nr3D [Achlioptas et al., 2020] dataset. 3) Appearance Captioning task from Chen et al. [2024c], where the model is asked to captioning the physical attributes or visual characteristics of a specified object. The ground truth is generated by CogVLM [Wang et al., 2023a]. 4) Region Captioning task from Chen et al. [2024c], where the model is asked to caption the region encircling a specified object. The ground truth is organized by ChatGPT [OpenAI, ]. 5) End to end 3D Visual Grounding from Nr3D dataset [Achlioptas et al., 2020], where the model is not provided ground truth candidates, different from the original setting in Nr3D. 6) End to end 3D Visual Grounding from Sr3D+ dataset [Achlioptas et al., 2020], where the model is also not provided ground truth candidates, different from the original setting in Sr3D+.
Figure 4: (Upper) Pipeline to generate our Diverse Instruction data by the in-context learning of ChatGPT. (Lower) The one-shot examples for ChatGPT to rephrase the instruction-following data.
@@
4.3 Data Summary
In summary, our Robust Instruction Generation engine produces two types of data: 1) Adversarial Instruction data: a total of 344K samples, consisting of mixed positive and negative pairs, formulated into four new tasks. 2) Diverse Instruction data: a total of 508K samples, covering multiple tasks and various language styles, diversified through in-context learning from ChatGPT.
5 Experiments
5.1 Benchmarks and Metrics
We provide quantitative results on five widely-used 3D multimodal learning benchmarks: ScanRefer [Chen et al., 2020] for 3D Visual Grounding, Multi3DRefer [Zhang et al., 2023] for General 3D Visual Grounding including zero, single and multiple target objects, Scan2Cap [Chen et al., 2021] for 3D Dense Captioning, ScanQA [Azuma et al., 2022] for 3D Question Answering, and SQA3D [Ma et al., 2022] for 3D Situated Question Answering. The vision data are all based on the ScanNet dataset [Dai et al., 2017], which contains real world 3D point clouds across 1,513 indoor scenes with detailed object annotations. All these benchmarks follow the same data split as ScanNet.
We follow the standard evaluation metrics widely adopted in the respective benchmarks. For ScanRefer, we calculate accuracy at Intersection over Union (IoU) thresholds of 0.25 and 0.5 (Acc@0.25, Acc@0.5). For Multi3DRefer, we use the F1 score with IoU thresholds of 0.25 and 0.5 to measure performance. In Scan2Cap, we apply the CIDEr@0.5 and BLEU-4@0.5 (C@0.5, B-4@0.5) metrics, combining standard captioning metrics with the IoU metric. For ScanQA, the METEOR and ROUGE metrics, denoted as M and R, are employed. Lastly, SQA3D is assessed with exact match accuracy (EM) and its extended form, EM-R, as suggested by LEO [Huang et al., 2023b].
5.2 Implementation Details
We extract 150 object features from each 3D scene, along with the corresponding position embeddings and 3D masks generated by Mask3D. The 2D Projector, as shown in Fig. 2, is a two-layer MLP. We use the Vicuna-7B-v1.5 model [Chiang et al., 2023] as our LLM and fine-tune it using LoRA [Hu et al., 2021] (with a rank of 16) by Cross Entropy loss. The global learning rate is formulated as [batch size $\times$ base learning rate $\times$ number of GPUs] and is set to 0.00064, with a cosine annealing schedule. For our results in Tab. 1, we first train for 2 epochs on the RIG-generated data, and then train for 2 epochs on the benchmark training sets in the second stage. We train for 1 epoch for each stage to efficiently conduct ablation studies of RIG-generated data. For ablation on RAP and IFB, we train for 1 epoch on the benchmark training sets to avoid potential compound effects.
@@ 5.4 Ablation Study
Table 2: Ablation study on Robust Instruction Generation. Benchmark denotes training on the original training set of the benchmarks. Adversarial denotes adding the Adversarial Instruction data to the training set. Diverse denotes adding the Diverse Instruction data to the training set.
@@
+ Adversarial 49.0 55.1 82.5 18.2 50.7
+ Diverse 50.6 53.1 80.5 18.0 50.9
+ Adversarial & Diverse 51.8 56.9 84.1 18.4 52.6
Table 3: Ablation study on our proposed modules: Relation-Augmented Projecter (RAP) and ID-Feature Bonding (IFB).
Model ScanRefer Multi3DRefer Scan2Cap ScanQA(val) SQA3D(val)
Acc@0.5 F1@0.5 C@0.5 M EM
Baseline 41.0 45.3 69.2 17.4 48.2
+ RAP 42.5 46.3 72.0 17.6 48.6
+ RAP & IFB 45.3 50.2 73.6 17.7 48.9
We perform ablation studies from the perspective of training data and model structure, respectively. We first evaluate the effectiveness of RIG-generated data by progressively adding the Adversarial Instruction data and the Diverse Instruction data to the training set. We then investigate the contribution of RAP and IFB by comparing models with and without these components.
Robust Instruction Generation (RIG): As shown in Tab. 2, by adding the Adversarial Instruction data, we observe a consistent improvement across all benchmarks. Specifically, performance on ScanRefer and Multi3DRefer increases by 3.7% and 4.9%, respectively. It is worth noting that the performance on Scan2Cap improves by 8.9%, even though there is not any object captioning data in the Adversarial Instruction data, which highlights its contribution on enhancing the understanding towards each object by challenging the model with mixed positive and negative samples. Additionally, by adding the Diverse Instruction data, we also provide comprehensive improvements. Specifically, descriptions from the original ScanRefer are annotated by human following a fixed instruction template or expression style, which limits the language diversity. In contrast, the Diverse Instruction data contains various language styles and task formats, which helps the model generalize better, resulting in a 5.3% improvement on ScanRefer. Finally, by combining both two types of data, we achieve a further improvement, demonstrating the effectiveness of RIG-generated data.
Relation-Augmented Projecter (RAP) and ID-Feature Bonding (IFB): As shown in Tab. 3, the integration of RAP leads to steady improvements across all benchmarks. Notably, the performance on Visual Grounding tasks, including ScanRefer and Multi3DRefer, shows significant gains due to RAP’s enhanced spatial comprehension. When IFB is incorporated, further improvements are observed, emphasizing the importance of refining the model’s object referring and grounding capabilities by reinforcing the connection between IDs and features.
5.5 Qualitative Results
Figure 5: Visualization of Robin3D’s responses on all the five benchmarks.
We provide visualization of Robin3D’s responses on all the benchmarks in Fig. 5 with the prompts of each task. These results demonstrates the generalization ability of Robin3D on various tasks.
6 Conclusion
To build a general-purpose AI agent in the 3D real world, we identify the problem of a lack of robust instruction training data in current 3DLLMs. To tackle this challenge, we introduce Robin3D, a powerful 3DLLM trained on large-scale instruction-following data generated by our novel data engine, Robust Instruction Generation (RIG) engine. We generate and collect 1 million instruction data, including benchmark data, adversarial data, and diverse data. To better handle these complex instructions, Robin3D incorporates a Relation-Augmented Projector to enhance the understanding of spatial relationships among objects, and ID-Feature Bonding for better object referring and grounding. Finally, Robin3D achieves state-of-the-art performance across five widely-used 3D multimodal learning benchmarks, making significant progress towards Spatial Intelligence.

### v2

@@ 3.2.2 Diverse Data Generation
We first collect large scale data from different tasks outside the benchmark dataset. Specifically, given a 3D scene, we collect question-answering pairs from the following tasks: 1) Category Question-Answering task from [20], where the model is asked to answer the category of a specified object. 2) Nr3D Captioning task from [20], where the model is asked to caption the spatial relationship of a specified object to its neighboor. The ground truth is constructed from Nr3D [1] dataset. 3) Appearance Captioning task from [9], where the model is asked to captioning the physical attributes or visual characteristics of a specified object. The ground truth is generated by CogVLM [43]. 4) Region Captioning task from [9], where the model is asked to caption the region encircling a specified object. The ground truth is organized by ChatGPT [37]. 5) End to end 3D Visual Grounding from Nr3D dataset [1], where the model is not provided ground truth candidates, different from the original setting in Nr3D. 6) End to end 3D Visual Grounding from Sr3D+ dataset [1], where the model is also not provided ground truth candidates, different from the original setting in Sr3D+.
@@
Figure 5: Overview of Robin3D model structure. Bottom: Relation-Augmented Projector fuses the features and position embedding from Mask3D and Uni3D for final 3D features. 2D features from DINO v2 are projected into the LLM space. Middle: Sequence Organization enhances the connection between object IDs and object features by wrapping the features with identical IDs and the Post-Vision order. Top: We use LoRA to fine-tune the LLM on our constructed 1 million instruction data.
3.3 Improvements on the model structure
We adopt Chat-Scene [19] as the baseline of our model structure, and further provide improvements on the vision projector and the organization of the sequence of tokens. Detailed and formal demonstration of our model structure is put in the Supplementary to avoid redundancy. We illustrate our improvements in the followings.
Vision Projector
Chat-Scene’s [19] 3D features inevitably weaken the spatial relationships between objects due to the object-centric normalization in Uni3D [50], which hinders the learning on our diverse visual grounding data. As shown in Fig. 5 (bottom), to obtain relation-aware 3D features while preserving the unified object-centric characteristics, our Relation-Augmented Projector (RAP) considers three types of 3D features: a) the object features from Mask3D, $X_{mask3d}$, which are scene-level respresentations, containing spatial relationship, as they come across multiple cross-attention layers to exchange information. b) the position embeddings of the Mask3D, $X_{pos}$, which are directly projected from the objects’ coordinates. c) the unified object features, $X_{uni3d}$, from Uni3D. Finally, our RAP is formulated as:
$\begin{split}X=\text{Concat}(&\text{Norm}_{L2}(X_{uni3d}),\text{Norm}_{L2}(X_{mask3d})),\\ X_{rap}&=\text{MLP}(X)+\text{MLP}(X_{pos})\end{split}$ (1)
where $\text{Norm}_{L2}$ is the L2 normalization, Concat is the concatenation alongside the channel dimension, and MLP is a multi-layer perceptron with GELU activation [16]. $X_{rap}$ represents our final 3D features.
Sequence Organization
Previous approaches [20, 19] simply append object ID to object feature as a prefix, which may lack sufficient connection between the ID and the feature, making it struggle with complex referring and grounding requirements in our adversarial instruction data. Here, we provide a more informative way to organize the input sequence. As shown in Fig. 5 (middle), we first use two identical ID tokens to wrap the object features. Adhering to the causal attention nature of the LLM, this approach links ID information to the object features via the first ID token, and links object information to its ID via the second ID token. Secondly, we adopt a post-vision order, which places the vision tokens at the end of the input sequence, closer to the answer tokens generated by the model. This approach mitigates attention deviation from the answer tokens to the ID-Feature pairs, a problem caused by their relative token distance and rotary position embeddings [41, 35]. The post-vision order is structured as: [$<$Question tokens$>$, $<$Vision tokens$>$, $<$Answer tokens$>$], where $<$Vision tokens$>$ comprises the ID tokens and object feature tokens.
4 Experiments
In this section, we present the main quantitative results and ablation studies. We have also conducted a comprehensive analysis of our Robust Instruction Data, along with more detailed ablation studies. However, due to space limitations, these are provided in the Supplementary. We highly recommend that readers refer to it for a complete understanding.
Figure 6: Ablation study on Robust Instruction Generation. Benchmark denotes training on the original training set of the benchmarks. Adversarial denotes adding the Adversarial Instruction data to the training set. Diverse denotes adding the Diverse Instruction data to the training set.
Figure 7: Ablation study on our proposed improvements: Relation-Augmented Projector and Sequence Organization.
4.1 Benchmarks and Metrics
We provide quantitative results on the widely-used 3D multimodal learning datasets: ScanRefer [6] for 3D Visual Grounding, Multi3DRefer [49] for General 3D Visual Grounding including zero, single and multiple target objects, Scan2Cap [10] for 3D Dense Captioning, ScanQA [3] for 3D Question Answering, and SQA3D [36] for 3D Situated Question Answering. The vision data are all based on the ScanNet dataset [12], which contains real world 3D point clouds across 1,513 indoor scenes with detailed object annotations. All these benchmarks follow the same data split as ScanNet.
We follow the standard evaluation metrics widely adopted in the respective benchmarks. For ScanRefer, we calculate accuracy at Intersection over Union (IoU) thresholds of 0.25 and 0.5 (Acc@0.25, Acc@0.5). For Multi3DRefer, we use the F1 score with IoU thresholds of 0.25 and 0.5 to measure performance. In Scan2Cap, we apply the CIDEr@0.5 and BLEU-4@0.5 (C@0.5, B-4@0.5) metrics, combining standard captioning metrics with the IoU metric. For ScanQA, the METEOR and ROUGE metrics, denoted as M and R, are employed. Lastly, SQA3D is assessed with exact match accuracy (EM) and its extended form, EM-R, as suggested by LEO [21].
4.2 Implementation Details
We extract 150 object features from each 3D scene, along with the corresponding position embeddings and 3D masks generated by Mask3D. Following Chat-Scene’s setup, we use a two-layer MLP as the 2D Projector and the Vicuna-7B-v1.5 model [11] as our LLM. We fine-tune the model using LoRA [18] (with a rank of 16) by Cross Entropy loss. The global learning rate is formulated as [batch size $\times$ base learning rate $\times$ number of GPUs] and is set to 0.00064, with a cosine annealing schedule. For our results in Tab. 1, we train 2 epochs on the RIG-generated data, and then train 2 epochs on the benchmark data in the second stage. We train 1 epoch for each stage to efficiently conduct ablation studies of RIG-generated data. For ablation on our improvements of the model structure, we train 1 epoch on the benchmark data to avoid potential compound effects.
Figure 8: Visualization of Robin3D’s responses on all the benchmarks.
@@ 4.4 Ablation Study
We perform ablation studies on the training data and model structure, respectively. We first evaluate the effectiveness of RIG-generated data by progressively adding the Adversarial Instruction data and the Diverse Instruction data to the training set. We then investigate the contribution of Relation-Augmented Projector and Sequence Organization by comparing models with and without these components.
Robust Instruction Generation (RIG): As shown in Fig. 6, by adding the Adversarial Instruction data, we observe a consistent improvement across all benchmarks. Specifically, performance on ScanRefer and Multi3DRefer increases by 3.7% and 4.9%, respectively. It is worth noting that the performance on Scan2Cap improves by 8.9%, even though there is not any object captioning data in the Adversarial Instruction data, which highlights its contribution on enhancing the understanding towards each object by challenging the model with mixed positive and negative samples. Additionally, by adding the Diverse Instruction data, we also provide comprehensive improvements. Specifically, descriptions from the original ScanRefer are annotated by human following a fixed instruction template or expression style, which limits the language diversity. In contrast, the Diverse Instruction data contains various language styles and task formats, which helps the model generalize better, resulting in a 5.3% improvement on ScanRefer. Finally, by combining both two types of data, we achieve a further improvement, demonstrating the effectiveness of RIG-generated data.
Relation-Augmented Projector and Sequence Organization: As shown in Fig. 7, the baseline indicates the Chat-Scene’s structure and the integration of our projector leads to steady improvements across all benchmarks. Notably, the performance on Visual Grounding tasks, including ScanRefer and Multi3DRefer, shows significant gains due to our enhanced spatial comprehension. When adopting our sequence organization, further improvements are observed, emphasizing the importance of refining the model’s object referring and grounding capabilities by reinforcing the connection between IDs and features.
4.5 Qualitative Results
We provide the visualization of Robin3D’s responses on all the benchmarks in Fig. 8 with the prompts of each task. These results demonstrate the generalization ability of Robin3D on various tasks.
5 Conclusion
We identify the problem of a lack of robust instruction training data in current 3DLLMs. To tackle this challenge, we introduce Robin3D, a powerful 3DLLM trained on large-scale instruction-following data generated by our novel data engine, Robust Instruction Generation (RIG) engine. We generate and collect 1 million instruction data, including benchmark data, adversarial data, and diverse data. To better handle our complex data, Robin3D incorporates a Relation-Augmented Projector and enhanced sequence organization for better object referring and grounding. Finally, Robin3D achieves state-of-the-art performance across all the widely-used 3D multimodal learning benchmarks.
6 Future Work
Our main contribution lies in the novel approach of constructing robust data to enhance the model’s capabilities. We hope this work will inspire the current community to explore ways to further improve the complexity and quality of training data, rather than simply relying on available benchmark datasets. At the same time, developing a rigorous and comprehensive benchmark to better evaluate model performance is also a potentially valuable direction.
Additionally, our study follows previous common setting in 3DLLM [17, 7, 20, 19] by training and testing on the ScanNet indoor environment. Expanding future work to outdoor environments or to data based on LiDAR could further close the gap in 3DLLM and Spatial Intelligence.
It is worth noting that open-vocabulary capabilities have not yet been formally evaluated or emphasized in previous 3DLLM research. Training a powerful vision backbone that simultaneously supports open-vocabulary capabilities, unified representation, and scene-level representation could foreseeably lead to significant improvements.
Supplementary Material
Figure 9: The number of samples for different tasks in our robust data and the visualization of their proportion of the total data.
Figure 10: The word cloud of our adversarial data.
Figure 11: The word cloud of our diverse data.
Figure 12: The average sentence length of different tasks in our robust data.
7 Data Analysis
As shown in Fig. 9, we provide detailed statistics on the number of samples for different tasks in our robust dataset, along with a qualitative result of their respective proportions in the total data. For our Diverse Instruction data, we split it into three parts based on the task categories for statistical purposes, which include Diverse Visual Grounding, Diverse Captioning, and Diverse Question Answering.
We further present the word cloud of our Adversarial Instruction data and Diverse Instruction data in Fig. 10 and Fig. 11, respectively. We exclude the words related to object IDs, as they pertain to the referring and grounding format rather than the actual data content.
In Fig. 12, we provide statistics on the average sentence length for each task in our robust dataset. Here, the sentence length is calculated as the number of words in the question prompt plus the number of words in the answer, excluding the count of object IDs.
8 Detailed Ablation Study of Adversarial data
To further evaluate the effectiveness of each task in our Adversarial Instruction data, we conduct detailed ablation studies on Hybrid Object Probing Evaluation data, Hybrid Referring Object Classification data, Partial Factual 3D Visual Grounding data, and Faithful 3D Question Answering data by adding them to the benchmark data in each experiment. As shown in Tab. 2, all four tasks in the Adversarial data contribute notable improvements compared with solely training on the benchmark data.
@@ 8 Detailed Ablation Study of Adversarial data
+ HOPE 45.8 52.5 76.1 17.8 50.1
+ HROC 47.7 53.0 78.9 18.0 50.3
+ PF-3DVG 45.7 51.0 77.2 17.9 49.6
+ 3DFQA 47.2 52.1 77.5 17.9 50.2
Table 2: Ablation study on Adversarial Instruction data. Benchmark denotes training on the original training set of the benchmarks. HOPE denotes adding the Hybrid Object Probing Evaluation data to the original training set. HROC denotes adding the Hybrid Referring Object Classification data to the original training set. PF-3DVG denotes adding the Partial Factual 3D Visual Grounding data to the original training set. 3DFQA denotes adding the Faithful 3D Question Answering data to the original training set.
Figure 13: Overview of Robin3D model structure. Bottom: Our Relation-Augmented Projecter fuses the features and position embedding from Mask3D and Uni3D to generate final 3D features. 2D features from DINO v2 are projected into the LLM space. We freeze the Mask3D, Uni3D, and DINO v2. Middle: We enhance the connection between object IDs and object features by wrapping the features with identical IDs and the Post-Vision order. Top: We use LoRA to fine-tune the LLM on our constructed 1 million instruction data.
9 Details of Robin3D
To train a 3D LLM using instruction fine-tuning, we first represent the 3D scene as a sequence of vision tokens, then append it with system and instruction prompts, expressed as sequences of language tokens, to indicate the task. Taking the above tokens as input, a LLM is supervised to output the answer tokens via next token prediction. Specifically, as shown in Fig. 13, given the point cloud of a 3D scene, we use the pre-trained 3D segmenter Mask3D [40] to extract object features along with their corresponding 3D masks. Following Chat-Scene [19], we further sample each object’s point cloud based on the 3D masks, normalize it, and employ the pre-trained Uni3D [50] to extract unified object-centric 3D features. Additionally, 2D masks projected from the 3D masks are used to sample and average 2D features, which are extracted by DINO v2 from multi-view images of each object. Our Relation-Augmented Projector fuses the 3D features and position embeddings from Mask3D and Uni3D into our final 3D features. In line with Chat-Scene [19], we incorporate special tokens $\{<\texttt{OBJ}_{i}>\}_{i=1...n}$ as object IDs into the vocabulary. These ID tokens are paired with 2D and 3D object features to indicate each object, for referring to the object in the input instruction or grounding the object in model’s output. We combine each object feature with its corresponding object ID, and appends the system and question prompts at the beginning of the sequence, which are then fed into the LLM.